OpenStudy (anonymous):

given that z=(16-q)/(7-p) find the possible values of p and q that has: a) no solutions b) infinite solutions c) a unique solution

OpenStudy (anonymous):

a) p,q,z are elements of N b) p,q,z are elements of R c) p,q,z are alements of Q (not sure if that is right)

Directrix (directrix):

z=(16-q)/(7-p) -------------- a) If p = 7 and q is not 16, the denominator of the rational expression becomes 0, and there is no solution. b) If p = 7 and q = 16, the fraction becomes 0/0 which is indeterminate. c) If p is not 7, and if p and q are Real Numbers, I think there will be a unique solution. I am not certain that these thoughts are correct. Someone else will come along and add other ideas, I hope. @UnkleRhaukus

OpenStudy (anonymous):

@Directrix you are right~

Directrix (directrix):

@Kystal --> Thanks for letting me know.

OpenStudy (anonymous):

Do you know why 0/0 is indeterminate? I thought 0/0 gives 0?

Directrix (directrix):

 0/0 does equal 0 but it could equal any other real number, too. Think about how division is defined in terms of multiplication. 6 divided by 2 = 3 if 3*2 =6. What happens with 0/0, I'll draw. Does that make sense?